# Epstein Surfaces, $W$-Volume, and the Osgood-Stowe Differential

### v1

@@ 6.1 Univalence and quasidisk criteria after Epstein
Let $\hat{g}$ be a complete metric on $\Delta$, and $-^f\colon\Delta\to\widehat{{\mathbb{C}}}$ a locally univalent map such that
@@
Let $\hat{g}$ be a complete metric on $\Delta$, $-^f\colon\Delta\to\widehat{{\mathbb{C}}}$ a locally univalent map and $\hat{B}=B(\Sigma_{f},\hat{g})$ be the projective shape operator.
@@
The are two approaches to the proof of this theorem. In one approach in involves showing that the embedding $f$ extends continuously to the boundary the boundary of $\Delta$. This can be done be considering the limiting behavior of curves in ${{\mathbb{H}}^{3}}$ of curvature $\leq 1$. Instead, we follow a more classical approach using Carathéodory’s theory of prime ends which we now briefly describe.
@@ 6.1.1 Extension of univalent maps, prime ends
Let $-^f:\Delta\rightarrow U\subseteq\hat{\mathbb{C}}$ be a univalent map. Then $f$ extends continuously to $\partial\Delta$ if and only if $\partial U$ is locally connected.
@@ 6.1.2 Continuous extension and quasidisk criterion
Let $\hat{g}$ be a complete metric on $\Delta$, and $-^f\colon\Delta\to\widehat{{\mathbb{C}}}$ a locally univalent map such that projective shape operator $\hat{B}=B(\Sigma_{f},\hat{g})$ has non-negative eigenvalues. Then the map $f$ is univalent and $f$ extends continuously to its boundary.
@@
Let $-^K\subseteq\hat{\mathbb{C}}$ be closed with complement two disjoint open simply connected sets $U,V$ such that $K$ is accessible from both $U$ and $V$. Then $K$ is a Jordan curve.
@@
Let $\hat{g}$ be a complete metric on $\Delta$, and $-^f\colon\Delta\to\widehat{{\mathbb{C}}}$ a locally univalent map such that $\hat{B}=B(\Sigma_{f},\hat{g})$ have positive eigenvalues. Then $U$ is a Jordan domain and $f$ extends to a homeomorphism on $\overline{\Delta}$.
@@ 6.2 Quasi-conformal reflections
Let $\hat{g}$ be a complete metric on $\Delta$, $-^f\colon\Delta\to\widehat{{\mathbb{C}}}$ a locally univalent map and $\hat{B}=B(\Sigma_{f},\hat{g})$ be the projective shape operator. If $\hat{g}$ is negatively curved and there is a $10\leq k<$ 1 such that
@@
Proof: We keep the same notation in case 3 of Theorem 6.2 above. Then we let $H:\hat{\mathbb{C}}\rightarrow\hat{\mathbb{C}}$ be given by $H=f_{{\operatorname{ext}}}\circ C\circ f_{{\operatorname{ext}}}^{-1}$ where $C(z)=1/\overline{z}$. Trivially $H(z)=z$ on $\partial U$ and as one of either $f_{{\operatorname{ext}}}$ or $f_{{\operatorname{ext}}}^{-1}$ is conformal on each component, we have the beltrami differential bound is the same as for $f_{{\operatorname{ext}}}$. $\Box$
@@ 6.2.1 History
The study of these two problems has a long history. Two foundational results are the Nehari univalence criterion that if $\|Sf\|_{\infty}\leq\frac{1}{2}$ then $f$ is univalent (see [Neh]) and the Ahlfors-Weill Extension Theorem that if $\|Sf\|_{\infty}\leq\frac{k}{2}<\frac{1}{2}$, then $f$ extends to a $K=\frac{1+k}{1-k}$ quasiconformal homeomorphism (see [AW]). These results can be obtained by letting $\hat{g}$ be they hyperbolic metric in Theorems 6.1 and 6.2
@@
As well as generalizing Epstein’s criteria, our approach is short, proving the main results of [Eps2, Eps3, Eps4] and does not require the use of Epstein’s unpublished paper [Eps1] where he proved a version of Shur’s Lemma for hyperbolic space. Also as Epstein’s first paper [Eps4] (where he introduced Epstein surfaces and proved many of their properties) remains unpublished, we hope that this paper will bring the results in this important paper to a wider audience.
@@ 6.3 Example $f(z)=z^{c}$
Letting $h(\theta)=-\log(\sin(\theta))$ gives the hyperbolic metric and therefore we obtain Nehari’s condition. We now let $h(\theta)=-2\log(\sin(\theta))$. Then the above gives
@@
After some elementary algebra, this reduces to
@@
To compare, this gives a better univalence condition than the Nehari inequality $||S(f)||_{\infty}<1/2$ which is $|c^{2}-1|\leq 1$ and is a closer approximation to the full set of univalent maps of the form $z^{c}$ which is the set $|c-1|\leq 1$ (see Figure 1).
Figure 1: Comparing univalence conditions for $f(z)=z^{c}$
If $-1$ is an eigenvalue of $\hat{B}$ then it is still possible to define the triple $(f,g,B)$ however, where there are eigenvalues of $-1$ the map $f$ will be singular. To do this we note that if $\hat{\lambda}$ is an eigenvalue of $\hat{B}$ then $e^{-2t}\hat{\lambda}$ is an eigenvalue of $\hat{B}_{t}$. In particular, for all $p\in S$ there are at most two values $t$ where $-1$ is an eigenvalue of $\hat{B}_{t}$ so $f_{t}$ and $\hat{f}_{t}$ are defined for all but at most two $t$. If $-1$ is an eigenvalue of $\hat{B}_{t_{0}}$ and is not an eigenvalue of $\hat{B}_{t_{1}}$ then we define $\hat{f}_{t_{0}}(p)={\mathfrak{g}}_{t_{0}-t_{1}}\circ\hat{f}_{t_{1}}(p)$.margin: Orphan Black
If $-1$ is not an eigenvalue of $\hat{B}$ then by Theorem 5.1, the dual pair $(g,B)$ satisfies the Gauss-Codazzi equations in ${{\mathbb{H}}^{3}}$. If $S$ is simply connected then by Theorem 5.12, there is an immersion
$f\colon S\to{{\mathbb{H}}^{3}}$
with $g=f^{*}g_{{\mathbb{H}}^{3}}$ and shape operator $B$. Note that $-1$ will not be an eigenvalue of $B$ so by Lemma 5.14, the composition $f_{\infty}={\mathfrak{g}}_{\infty}\circ\hat{f}$ is an immersion.
@@ 7.2 Convexity
Let $(\Sigma,\hat{g},\hat{B})$ be a projective structure, conformal metric and shape operator with dual triple $(f,g,B)$. Let $f_{t}=\pi\circ{\mathfrak{g}}_{t}\circ\hat{f}$. Then for
@@
Proof: The map $f_{t}$ has shape operator $B_{t}=({\operatorname{Id}}+e^{-2t}\hat{B})^{-1}({\operatorname{Id}}-e^{-2t}\hat{B})$. Therefore if $\lambda_{0},\lambda_{1}$ are the eigenvalues of $\hat{B}$ the surface given by $f_{t}$ has principal curvatures
$\frac{1-e^{-2t}\lambda_{i}}{1+e^{-2t}\lambda_{i}}.$
If $e^{2t_{0}}>\max\{|\lambda_{0}|,|\lambda_{1}|\}$ these are both positive. Therefore if this holds at every point then $f_{t_{0}}$ is an immersion and locally convex. Therefore by the prior lemma
$\max\{|\lambda_{0}|,|\lambda_{1}|\}=|K(\hat{g})|+4\|Q(\Sigma,\hat{g})\|.$
Therefore $f_{t}$ is locally convex for $t>\sup\{|K(\hat{g})|+4\|Q(\Sigma,\hat{g})\|\}$. By [BBB], if a complete conformal metric $\hat{g}$ on $\Sigma$ has locally convex immersed surface, then $g_{\Sigma}\leq\hat{g}$ where $g_{\Sigma}$ is the projective metric on $\Sigma$. Therefore for $t>\sup\{|K(\hat{g})|+4\|Q(\Sigma,\hat{g})\|\}$ we have $g_{\Sigma}\leq e^{2t}\hat{g}.$ Taking the infimum over all such $t$ we get
@@ 9.1 Conformal variations
where normal-$\nabla$ is the gradient of $\hat{g}$.
@@ 9.3 $W(\hat{g}_{h},\hat{g}_{\Sigma})$
To prove Theorem 2.10, we will obtain upper and lower bounds on $W(\hat{g}_{h},\hat{g}_{\Sigma})$. As the projective metric $\hat{g}_{\Sigma}$ is not smooth, we first need to extend $W$. We define ${\cal G}^{0}_{c}(\Sigma)=\overline{{\cal G}^{0}_{c}(\Sigma)}\subseteq{\cal G}^{0}(\Sigma)$ where the closure is with respect to the sup-norm topology on the space of continuous conformal metrics ${\cal G}^{0}(\Sigma)$. In the appendix we will show that $W$ extends continuously to ${\cal G}^{0}_{c}(\Sigma)\times{\cal G}^{0}_{c}(\Sigma)$ and that $\hat{g}_{\Sigma}\in{\cal G}^{0}_{c}(\Sigma)$. This will suffice for our purposes.
@@ A.1 Hyperbolic manifolds
As we noted above a convex submanifold $N$ of a convex co-compact hyperbolic 3-manifold $M$ determines a conformal metric $\hat{g}$ on $\partial\bar{M}$. We would like to define the $W$-volume for $\hat{g}$ (and $N$) even when $\hat{g}$ is not smooth. While the volume term (the volume of $N$) is still well defined, the integral of mean curvature is not.
For these non-smooth convex manifolds it will be convenient to fix the ambient convex co-compact hyperbolic 3-manifold $M$. In Lemma 8.3 we saw that convex submanifolds of $M$ determine conformal metrics on $\partial_{c}M$. We let ${\cal G}^{0}_{c}(\partial\bar{X};M)$ be the space of such metrics. This will contain the space of smooth metrics denoted ${\cal G}_{c}(\partial\bar{X};M)$. As all metrics are conformally equivalent given $\hat{g}$ and $\hat{h}$ in ${\cal G}^{0}_{c}(\partial\bar{X};M)$ we have $\hat{h}=e^{2u}\hat{g}$ for some continuous function $u$. We then define the sup-norm by $\|\hat{g}-\hat{h}\|_{\infty}=\|u\|_{\infty}$. We also have $\hat{h}\geq\hat{g}$ if $u\geq 0$ and $\hat{h}\leq\hat{g}$ if $u\leq 0$.
@@
For this result to be useful we need to know that smooth metrics are dense in ${\cal G}_{c}^{0}(\partial\bar{X};M)$.

### v2

@@ 6.1 Univalence and quasidisk criteria after Epstein
Let $\hat{g}$ be a complete metric on $\Delta$, and $f\colon\Delta\to\widehat{{\mathbb{C}}}$ a locally univalent map such that
@@
Let $\hat{g}$ be a complete metric on $\Delta$, $f\colon\Delta\to\widehat{{\mathbb{C}}}$ a locally univalent map and $\hat{B}=B(\Sigma_{f},\hat{g})$ be the projective shape operator.
@@
The are a number of natural approaches to the proof of this theorem. One involves showing that the embedding $f$ extends continuously to the boundary by considering the limiting behavior of curves in ${{\mathbb{H}}^{3}}$ of curvature $\leq 1$. Instead, we follow a more classical approach using Carathéodory’s theory of prime ends which we now briefly describe.
@@ 6.1.1 Extension of univalent maps, prime ends
Let $f:\Delta\rightarrow U\subseteq\hat{\mathbb{C}}$ be a univalent map. Then $f$ extends continuously to $\partial\Delta$ if and only if $\partial U$ is locally connected.
@@ 6.1.2 Continuous extension and quasidisk criterion
Let $\hat{g}$ be a complete metric on $\Delta$, and $f\colon\Delta\to\widehat{{\mathbb{C}}}$ a locally univalent map such that projective shape operator $\hat{B}=B(\Sigma_{f},\hat{g})$ has non-negative eigenvalues. Then the map $f$ is univalent and $f$ extends continuously to its boundary.
@@
Let $K\subseteq\hat{\mathbb{C}}$ be closed with complement two disjoint open simply connected sets $U,V$ such that $K$ is accessible from both $U$ and $V$. Then $K$ is a Jordan curve.
@@
Let $\hat{g}$ be a complete metric on $\Delta$, and $f\colon\Delta\to\widehat{{\mathbb{C}}}$ a locally univalent map such that $\hat{B}=B(\Sigma_{f},\hat{g})$ have positive eigenvalues. Then $U$ is a Jordan domain and $f$ extends to a homeomorphism on $\overline{\Delta}$.
@@ 6.2 Quasi-conformal reflections
Let $\hat{g}$ be a complete metric on $\Delta$, $f\colon\Delta\to\widehat{{\mathbb{C}}}$ a locally univalent map and $\hat{B}=B(\Sigma_{f},\hat{g})$ be the projective shape operator. If $\hat{g}$ is negatively curved and there is a $10\leq k<$ 1 such that
@@
Proof: We keep the same notation in case 3 of Theorem 6.2 above. Then we let $H:\hat{\mathbb{C}}\rightarrow\hat{\mathbb{C}}$ be given by $H=f_{{\operatorname{ext}}}\circ C\circ f_{{\operatorname{ext}}}^{-1}$ where $C(z)=1/\overline{z}$. Trivially $H(z)=z$ on $\partial U$ and as one of either $f_{{\operatorname{ext}}}$ or $f_{{\operatorname{ext}}}^{-1}$ is conformal in the composition at any point in the complement of $\partial U$, we have the beltrami differential bound is the same as for $f_{{\operatorname{ext}}}$. $\Box$
@@ 6.2.1 History
The study of these two problems has a long history. Two foundational results are the Nehari univalence criterion that if $\|Sf\|_{\infty}\leq\frac{1}{2}$ then $f$ is univalent (see [Neh]) and the Ahlfors-Weill Extension Theorem that if $\|Sf\|_{\infty}\leq\frac{k}{2}<\frac{1}{2}$, then $f$ extends to a $K=\frac{1+k}{1-k}$ quasiconformal homeomorphism (see [AW]). These results can be obtained by letting $\hat{g}$ be they hyperbolic metric in Theorems 6.1 and 6.2.
@@
As well as generalizing Epstein’s criteria, our approach is short (sections 4 to 6 of this paper), proving the main results of [Eps2, Eps3, Eps4] and does not require the use of Epstein’s unpublished paper [Eps1] where he proved a version of Shur’s Lemma for hyperbolic space. Also as Epstein’s first paper [Eps4] (where he introduced Epstein surfaces and proved many of their properties) remains unpublished, we hope that this paper will bring the results in this important paper to a wider audience.
@@ 6.3 Example $f(z)=z^{c}$
Letting $h(\theta)=-\log(\sin(\theta))$ gives the hyperbolic metric and therefore we obtain Nehari’s condition $||S(f)||_{\infty}<1/2$ or equivalently $|c^{2}-1|\leq 1$.
We now let $h(\theta)=-2\log(\sin(\theta))$. Then we have
@@
After simplification, this reduces to
@@
To compare, this gives a better univalence condition than the Nehari inequality $|c^{2}-1|\leq 1$ and is a closer approximation to the full set of univalent maps of the form $z^{c}$ which is the set $|c-1|\leq 1$ (see Figure 1).
@@ 7.2 Convexity
Let $(\hat{g},\hat{B})$ be dual to $(g,B)$ with $\hat{B}=B(\Sigma,\hat{g})$. Let $f$ be an immersion into ${{\mathbb{H}}^{3}}$ with fundamental pair $(g,B)$ and $f_{t}=\pi\circ{\mathfrak{g}}_{t}\circ\hat{f}$. Then for
@@
Proof: As $\hat{B}_{t}=e^{-2t}\hat{B}$ and $\hat{B}=B(\Sigma,\hat{g})$ has eigenvalues $-K(\hat{g})\pm 4\|Q(\Sigma,\hat{g})\|$ then for
$e^{2t}>\sup\{|K(\hat{g})|+4\|Q(\Sigma,\hat{g})\|\}$
we have $\hat{B}_{t}$ has eigenvalues in $(-1,1)$. Thus by duality, the eigenvalues of $B$ are in $(0,\infty)$. It follows that $f_{t}$ has positive principal curvatures and therefore is a locally convex immersion.
By [BBB], if $\hat{g}$ is complete and pair $(\hat{g},\hat{B})$ has locally convex immersed dual surface, then $g_{\Sigma}\leq\hat{g}$ where $g_{\Sigma}$ is the projective metric on $\Sigma$. Therefore for $t>\sup\{|K(\hat{g})|+4\|Q(\Sigma,\hat{g})\|\}$ we have $g_{\Sigma}\leq e^{2t}\hat{g}.$ Taking the infimum over all such $t$ we get
@@ 9.1 Conformal variations
where $\nabla$ is the gradient of $\hat{g}$.
@@ 9.3 $W(\hat{g}_{h},\hat{g}_{\Sigma})$
To prove Theorem 2.10, we will obtain upper and lower bounds on $W(\hat{g}_{h},\hat{g}_{\Sigma})$. As the projective metric $\hat{g}_{\Sigma}$ is not smooth, we first need to extend $W$. We define ${\cal G}^{0}_{c}(\Sigma)=\overline{{\cal G}_{c}(\Sigma)}\subseteq{\cal G}^{0}(\Sigma)$ where the closure is with respect to the sup-norm topology on the space of continuous conformal metrics ${\cal G}^{0}(\Sigma)$. In the appendix we will show that $W$ extends continuously to ${\cal G}^{0}_{c}(\Sigma)\times{\cal G}^{0}_{c}(\Sigma)$ and that $\hat{g}_{\Sigma}\in{\cal G}^{0}_{c}(\Sigma)$. This will suffice for our purposes.
@@ A.1 Hyperbolic manifolds
As we noted above, a convex submanifold $N$ of a convex co-compact hyperbolic 3-manifold $M$ determines a conformal metric $\hat{g}$ on $\partial\bar{M}$. We would like to define the W-volume for $\hat{g}$ (and $N$) even when $\hat{g}$ is not smooth. While the volume term (the volume of $N$) is still well defined, the integral of mean curvature is not.
For these non-smooth convex manifolds it will be convenient to fix the ambient convex co-compact hyperbolic 3-manifold $M$. In Lemma 8.3 we saw that convex submanifolds of $M$ determine conformal metrics on $\partial_{c}M$. We let ${\cal G}^{0}_{c}(\partial\bar{X};M)$ be the space of such metrics. This will contain the space of smooth metrics denoted ${\cal G}_{c}(\partial\bar{X};M)$. As all metrics are conformally equivalent, given $\hat{g}$ and $\hat{h}$ in ${\cal G}^{0}_{c}(\partial\bar{X};M)$, we have $\hat{h}=e^{2u}\hat{g}$ for some continuous function $u$. We then define the sup-norm by $\|\hat{g}-\hat{h}\|_{\infty}=\|u\|_{\infty}$. We also have $\hat{h}\geq\hat{g}$ if $u\geq 0$ and $\hat{h}\leq\hat{g}$ if $u\leq 0$.
@@
To extend our results for smooth metrics to continuous metrics we further need to know that smooth metrics are dense in ${\cal G}_{c}^{0}(\partial\bar{X};M)$.